Here is my problem. I have state the problem, as it has been put to us and just below it, I have shown what I have done to this point. I'm not sure I can get any more of it and make any sense with it. Can you help? Thanks in advance.
When using the quadratic formula to solve a quadratic equation ax2 + bx + c = 0, the discriminant is b2 - 4ac. This discriminant can be positive, zero, or negative. (When the discriminant is negative, then we have the square root of a negative number. This is called an imaginary number, sqrt(-1) = i. )
• Explain what the value of the discriminant means to the graph of y = ax2 + bx + c. Hint: Chose values of a, b and c to create a particular discriminant. Then, graph the corresponding equation.
• Search the Library and Internet. In the real world, where might these imaginary numbers be used?


ax2 + BX + C = 0


Move terms containing A to the left and the others to the right.


ax2 = - BX - C


ax2 = - (-XB+C)


Here Divide by X to isolate A.

-
A = + - (-XB+C)
2
X

ax2 + BX + C = 0


Solution


A = +- (-XB+C) ____
2
X

Answer by stanbon(75887)

• Explain what the value of the discriminant means to the graph of y = ax2 + bx + c. Hint: Chose values of a, b and c to create a particular discriminant. Then, graph the corresponding equation.
---------
Do that:
y=2x^2+3x+0 will pass thru the x axis twice because b^2-4ac>0
y=2x^2+3x+7 will not touch the x axis because b^2-4ac<0
y=x^2+4x+4 will touch but not pass thru the x axis because b^2-4ac=0
-------------------
• Search the Library and Internet. In the real world, where might these imaginary numbers be used?
Use Google: Imaginary numbers are used in modeling electricity, for example.
--------
Cheers,
Stan H.